Dry air will break down if the electric field exceeds about3.0 × 106 V/m. What amount of charge can be placedon a capacitor if the area of each plate is 3.1 cm2?

                  nC
I tried using this formula rearranged for Q :
Q = Eε0A
    = (3.0 × 106 V/m)(8.85 x10-12 Nm2/C2)(.031m2) = 8.2305 x 10-7C = 8.2405e-7nC
However, this is not right. Can anyone tell me what I amdoing wrong?
Answer #1
Concepts and reason

The concepts used to solve this problem are charge stored on the plates of capacitor and

Electric field for a parallel plate capacitor.

Initially, use the equation of electric field between the plates of capacitor, re-arrange the equation for Q, substitute the values and find the result.

Fundamentals

The electric field between parallel plates of a parallel plate capacitor is given as follows:

E=σε0E = \frac{\sigma }{{{\varepsilon _0}}}

Here, σ\sigma is the surface charge density and ε0{\varepsilon _0} is the electrical permittivity of vacuum.

The surface charge density is given as follows:

σ=QA\sigma = \frac{Q}{A}

Here, Q is the charge stored and A is the area.

Substitute equation σ=QA\sigma = \frac{Q}{A} in equation E=σε0E = \frac{\sigma }{{{\varepsilon _0}}} and re-arrange the expression for Q.

E=σε0=QAε0Q=AEε0\begin{array}{c}\\E = \frac{\sigma }{{{\varepsilon _0}}}\\\\ = \frac{Q}{{A{\varepsilon _0}}}\\\\Q = AE{\varepsilon _0}\\\end{array}

Substitute 3.0×106V/m3.0 \times {10^6}{\rm{ V/m}} for E, 3.1cm23.1{\rm{ c}}{{\rm{m}}^2} for A and 8.85×1012Nm2/C28.85 \times {10^{ - 12}}{\rm{ N}} \cdot {{\rm{m}}^2}/{{\rm{C}}^2} for ε0{\varepsilon _0} in equation Q=AEε0Q = AE{\varepsilon _0} .

Q=(3.1cm2(104m21.0cm2))(3.0×106V/m)(8.85×1012Nm2/C2)=8.23×109C=8.23nC\begin{array}{c}\\Q = \left( {3.1{\rm{ c}}{{\rm{m}}^2}\left( {\frac{{{{10}^{ - 4}}{\rm{ }}{{\rm{m}}^2}}}{{1.0{\rm{ c}}{{\rm{m}}^2}}}} \right)} \right)\left( {3.0 \times {{10}^6}{\rm{ V/m}}} \right)\left( {8.85 \times {{10}^{ - 12}}{\rm{ N}} \cdot {{\rm{m}}^2}/{{\rm{C}}^2}} \right)\\\\ = 8.23 \times {10^{ - 9}}{\rm{ C}}\\\\ = 8.23{\rm{ nC}}\\\end{array}

Ans:

The amount of charge stored is 8.23 nC.